# Clearer Frames, Anytime: Resolving Velocity Ambiguity in Video Frame Interpolation

### v2

@@ 5.3 Quantitative comparison
Table 2: Ablation study of the number of iterations on Vimeo90K septuplet dataset. $[\cdot]^{\#}$ denotes the number of iterations used for inference.
@@
In realistic scenarios where the precise distance map is inaccessible at inference, one could resort to a uniform map akin to time indexing. The bottom segment of Tab. 1 shows the performance of the enhanced models $[D]$ and $[D,R]$, utilizing identical inputs as model $[T]$. Given the misalignment between predicted frames using a uniform distance map and the ground-truth, the enhanced models do not outperform the base model on pixel-centric metrics. However, we argue that in most applications, the goal of VFI is not to predict pixel-wise aligned frames, but to generate plausible frames with high perceptual quality. Furthermore, pixel-centric metrics are less sensitive to blur [51], the major artifact introduced by velocity ambiguity. The pixel-centric metrics are thus less informative and denoted in gray. On perceptual metrics (especially NIQE), the enhanced models significantly outperforms the base model. This consistency with our qualitative observations further validates the effectiveness of distance indexing and iterative reference-based estimation.
Ablation study of the number of iterations.
Tab. 2 offers an ablation study on the number of iterations and the efficacy of a pure iterative reference-based estimation strategy. The upper section suggests that setting iterations at two strikes a good trade-off between computational efficiency and performance. The lower segment illustrates that while iterative reference-based estimation generally works for time indexing, there are exceptions, as observed with AMT-S. However, when combined with distance indexing, iterative reference-based estimation exhibits more stable improvement, as evidenced by the results for $[D,R]_{u}$. This is consistent with qualitative comparison.
Table 3: Comparison on Adobe240 [36] for $\times$8 interpolation with RIFE [11].
@@
Table 4: Comparison on X4K1000FPS [34] for $\times$8 interpolation with RIFE [11].
@@
Table 5: Comparison on X4K1000FPS [34] for $\times$16 interpolation with RIFE [11].
@@
Table 6: Comparison on Vimeo90K [47] using GMFlow [45] for distance map calculation with RIFE [11].
@@
Comparison on other benchmarks.
The septuplet set of Vimeo90K [47] is large enough to train a practical video frame interpolation model, and it represents the situations where the temporal distance between input frames is large. Thus, Vimeo90K (septuplet) can well demonstrate the velocity ambiguity problem that our work aims to highlight. Nonetheless, we report more results on other benchmarks. Tab. 6 and Tab. 6 show the results of RIFE [11] on Adobe240 [36] and X4K1000FPS [34] for $\times$8 interpolation respectively, using uniform maps. Distance indexing $[D]_{u}$ and iterative reference-based estimation $[R]_{u}$ strategies can consistently help improve the perceptual quality. In addition, it is noteworthy that $[D]_{u}$ is better than $[T]$ in terms of the pixel-centric metrics like PSNR, showing that the constant speed assumption (uniform distance maps) holds well on these two easier benchmarks. We further show $\times$16 interpolation on X4K1000FPS with larger temporal distance in Tab. 6. The results highlight that the benefits of our strategies are more pronounced with increased temporal distances.
Other optical flow estimator.
We also employ GMFlow [45] for the precomputation of distance maps, enabling an analysis of model performance when integrated with alternative optical flow estimations. The results are as shown in Tab. 6. Our strategies still lead to consistent improvement on perceptual metrics. However, this more recent and performant optical flow estimator does not introduce improvement compared to RAFT [38]. A likely explanation is that since we quantify the optical flow to $[0,1]$ scalar values for better generalization, our training strategies are less sensitive to the precision of the optical flow estimator.
@@
5.4 User study
To validate the effectiveness of our proposed strategies, we further conducted a user study with 30 anonymous participants. Participants were tasked with ranking the interpolation quality of frames produced by four model variants: [T], [D], [T,R], and [D,R]. See details of user study UI in supplementary materials. The results, presented in Fig. 7, align with our qualitative and quantitative findings. The [D,R] model variant emerged as the top-rated, underscoring the effectiveness of our strategies.
5.5 2D manipulation of frame interpolation
@@ 5.4 User study
6 Conclusion and Future Work
@@
We challenge the traditional time indexing paradigm and address its inherent uncertainties related to velocity distribution. Through the introduction of distance indexing and iterative reference-based estimation strategies, we offer a transformative paradigm to VFI. Our innovative plug-and-play strategies not only improves the performance in video interpolation but also empowers users with granular control over interpolation patterns across varied objects. Estimating accurate distance ratio maps from multiple frames represents a direction for our future research. Furthermore, the insights gleaned from our strategies have potential applications across a range of tasks that employ time indexing, such as space-time super-resolution, future predictions, blur interpolation and more.
Acknowledgement
@@
This work was supported by the Shanghai Artificial Intelligence Laboratory. We thank Dorian Chan, Zhirong Wu, and Stephen Lin for their insightful feedback and advice. Our thanks also go to Vu An Tran for developing the web application, and to Wei Wang for coordinating the user study.

### v3

@@ VI-C2 Comparison on Vimeo90K septuplet dataset
TABLE II: Ablation study of the number of iterations on Vimeo90K septuplet dataset. $[\cdot]^{\#}$ denotes the number of iterations used for inference.
@@
In realistic scenarios where the precise distance map is inaccessible at inference, one could resort to a uniform map akin to time indexing. The bottom segment of Table I shows the performance of the enhanced models $[D]$ and $[D,R]$, utilizing identical inputs as model $[T]$. Given the misalignment between predicted frames using a uniform distance map and the ground-truth, the enhanced models do not outperform the base model on pixel-centric metrics. However, we argue that in most applications, the goal of VFI is not to predict pixel-wise aligned frames, but to generate plausible frames with high perceptual quality. Furthermore, pixel-centric metrics are less sensitive to blur [56], the major artifact introduced by velocity ambiguity. The pixel-centric metrics are thus less informative and denoted in gray. On perceptual metrics (especially NIQE), the enhanced models significantly outperforms the base model. This consistency with our qualitative observations further validates the effectiveness of distance indexing and iterative reference-based estimation.
VI-C3 Ablation study of the number of iterations
Table II offers an ablation study on the number of iterations and the efficacy of a pure iterative reference-based estimation strategy. The upper section suggests that setting iterations at two strikes a good trade-off between computational efficiency and performance. The lower segment illustrates that while iterative reference-based estimation generally works for time indexing, there are exceptions, as observed with AMT-S. However, when combined with distance indexing, iterative reference-based estimation exhibits more stable improvement, as evidenced by the results for $[D,R]_{u}$. This is consistent with qualitative comparison.
TABLE III: Comparison on Adobe240 [58] for $\times$8 interpolation with RIFE [7].
@@ VI-C3 Ablation study of the number of iterations
TABLE IV: Comparison on X4K1000FPS [22] for $\times$8 interpolation with RIFE [7].
@@
TABLE V: Comparison on X4K1000FPS [22] for $\times$16 interpolation with RIFE [7].
@@
TABLE VI: Comparison on Vimeo90K [54] using GMFlow [59] for distance map calculation with RIFE [7].
@@
VI-C4 Comparison on other benchmarks
The septuplet set of Vimeo90K [54] is large enough to train a practical video frame interpolation model, and it represents the situations where the temporal distance between input frames is large. Thus, Vimeo90K (septuplet) can well demonstrate the velocity ambiguity problem that our work aims to highlight. Nonetheless, we report more results on other benchmarks. Table VI and Table VI show the results of RIFE [7] on Adobe240 [58] and X4K1000FPS [22] for $\times$8 interpolation respectively, using uniform maps. Distance indexing $[D]_{u}$ and iterative reference-based estimation $[R]_{u}$ strategies can consistently help improve the perceptual quality. In addition, it is noteworthy that $[D]_{u}$ is better than $[T]$ in terms of the pixel-centric metrics like PSNR, showing that the constant speed assumption (uniform distance maps) holds well on these two easier benchmarks. We further show $\times$16 interpolation on X4K1000FPS with larger temporal distance in Table VI. The results highlight that the benefits of our strategies are more pronounced with increased temporal distances.
VI-C5 Other optical flow estimator
We also employ GMFlow [59] for the precomputation of distance maps, enabling an analysis of model performance when integrated with alternative optical flow estimations. The results are as shown in Table VI. Our strategies still lead to consistent improvement on perceptual metrics. However, this more recent and performant optical flow estimator does not introduce improvement compared to RAFT [9]. A likely explanation is that since we quantify the optical flow to $[0,1]$ scalar values for better generalization, our training strategies are less sensitive to the precision of the optical flow estimator.
@@ VI-C5 Other optical flow estimator
VI-D User study
To validate the effectiveness of our proposed strategies, we further conducted a user study with 30 anonymous participants. Participants were tasked with ranking the interpolation quality of frames produced by four model variants: [T], [D], [T,R], and [D,R]. See details of user study UI in Appendix. The results, presented in Fig. 8, align with our qualitative and quantitative findings. The [D,R] model variant emerged as the top-rated, underscoring the effectiveness of our strategies.
VI-E 2D manipulation of frame interpolation
@@ VI-D User study
VI-F Multi-frame qualitative comparison
As shown in Fig. 10, we present a qualitative comparison of different variants of each VFI model under multi-frame setup. $[D,M]_{e}$ denotes the model that incorporate both the indexing map with multi-frame refiner and continuous indexing map estimator. As anticipated, the $[D,M]_{e}$ configuration consistently yields the highest visual quality across all base models, demonstrating the effectiveness of both the multi-frame refiner and the continuous indexing map estimator.
Figure 10: Qualitative comparison under multi-frame setting. $[T]$: original arbitrary time VFI models using time indexing. $[D]_{u}$: models trained using distance indexing, then inference using uniform maps. $[D,R]_{u}$: models trained using both distance indexing and iterative reference-based estimation, then inference using uniform maps. $[D,M]_{e}$: models trained using distance indexing with multi-frame fusion, then inference using estimated maps.
VI-G Multi-frame quantitative comparison
In Table VII, we present a quantitative performance comparison across various model variants. The configurations $[T]$, $[D]_{u}$, and $[D]$ retain the same definitions as previously described. $[D]_{e}$ denotes the model augmented with a continuous indexing map estimator, $[D,M]$ includes the proposed multi-frame fusion with the refiner model, and $[D,M]_{e}$ incorporates both the refiner and the map estimator. On pixel-level metrics such as PSNR and SSIM, the enhanced variant $[D,M]_{e}$ consistently outperforms the base model $[T]$ as well as $[D]_{u}$, which lacks a map estimator and relies on uniform indexing maps. The improvement from $[D]_{u}$ to $[D,M]_{e}$ highlights the effectiveness of integrating the continuous indexing map estimator. Meanwhile, performance gains from $[D]$ to $[D,M]$, and from $[D]_{e}$ to $[D,M]_{e}$, demonstrate the additional benefit provided by the multi-frame refiner.
TABLE VII: Multi-frame comparison on Vimeo90K septuplet dataset. $[M]$ denotes the multi-frame fusion. $[\cdot]_{e}$ denotes inference with estimated indexing map. Other notations have the same meaning as in the previous experiments.
RIFE [7] $[T]$ $[D]_{u}$ $[D]_{e}$ $[D,M]_{e}$ $[D]$ $[D,M]$
PSNR $\uparrow$ 28.22 27.55 28.25
28.34
29.20
31.63
SSIM $\uparrow$ 0.912 0.902 0.923
0.928
0.929
0.952
LPIPS $\downarrow$ 0.105 0.092 0.099
0.089
0.092
0.062
NIQE $\downarrow$ 6.663 6.344 6.554
6.173
6.475
5.990
IFRNet [24] $[T]$ $[D]_{u}$ $[D]_{e}$ $[D,M]_{e}$ $[D]$ $[D,M]$
PSNR $\uparrow$ 28.26 27.40 27.63
28.28
29.25
32.11
SSIM $\uparrow$ 0.915 0.902 0.907
0.919
0.931
0.958
LPIPS $\downarrow$ 0.088 0.083 0.087
0.083
0.080
0.074
NIQE $\downarrow$ 6.422
6.196
6.414 6.249 6.342
5.957
AMT-S [18] $[T]$ $[D]_{u}$ $[D]_{e}$ $[D,M]_{e}$ $[D]$ $[D,M]$
PSNR $\uparrow$ 28.52 27.33 27.60
28.80
29.61
31.80
SSIM $\uparrow$ 0.920 0.902 0.908
0.922
0.937
0.955
LPIPS $\downarrow$ 0.101 0.090 0.098
0.084
0.086
0.072
NIQE $\downarrow$ 6.866 6.382 6.452
6.223
6.656
6.056
EMA-VFI [17] $[T]$ $[D]_{u}$ $[D]_{e}$ $[D,M]_{e}$ $[D]$ $[D,M]$
PSNR $\uparrow$ 29.41 28.24 28.67
29.45
30.29
31.31
SSIM $\uparrow$ 0.928 0.912 0.919
0.932
0.942
0.951
LPIPS $\downarrow$ 0.086
0.079
0.082 0.084 0.078
0.069
NIQE $\downarrow$ 6.736 6.457 6.609
6.313
6.545
6.146
VI-H Multi-frame tuning strategy
In this section, we evaluate different training strategies for multi-frame fusion using RIFE [7] as the baseline for ablation studies. All experiments follow the same setting of previous experiments on Vimeo90k. We find that jointly tuning both refiner and map estimator leads to instability and failure to converge. Thus, we split the tuning process in two stages. In the first stage, we first finetune the multi-frame refiner $\mathcal{F^{\prime}}$ alone using original VFI loss. In the second stage, there are two possible choices. Since jointly tuning with the map estimator remains unstable, we either finetune the flow estimator only to adapt it to VFI (referred to as FE), or finetune the VFI modules $\mathcal{F}$ and $\mathcal{F^{\prime}}$ to adapt them to the pretrained flow estimator (referred to as FE), as shown in Table VIII. The loss term $\mathcal{L}_{D}$ is calculated by $\mathcal{L}_{D}=\|D_{t}-D_{t}^{\text{RAFT}}\|_{2}^{2}$, denoting the L2 loss between the indexing maps from CPFlow and RAFT. The setting $\mathcal{L}_{V_{f}}$ corresponds to training only the VFI refiner with its original loss, while $\mathcal{L}_{V_{b+f}}$ involves jointly training both the base VFI model and the refiner with their original loss. Among the tested strategies, jointly tuning both the VFI model and the refiner under $\mathcal{L}_{V_{b+f}}$ achieves the best performance. This improvement is attributed to the model’s ability to adapt both components to the updated indexing map derived from the newly predicted optical flow.
TABLE VIII: Multi-frame ablation study
FE VFI
$\mathcal{L}_{D}$ $\mathcal{L}_{V_{b+f}}$ $\mathcal{L}_{D+V_{b+f}}$ $\mathcal{L}_{V_{f}}$ $\mathcal{L}_{V_{b+f}}$
PSNR $\uparrow$ 26.10 27.22 27.34 30.89
31.63
SSIM $\uparrow$ 0.891 0.898 0.901 0.938
0.952
LPIPS $\downarrow$ 0.111 0.104 0.097 0.069
0.062
NIQE $\downarrow$ 6.627 6.575 6.509 6.087
5.990
TABLE IX: Costs comparison including runtime and number of parameters. We evaluate the computational overhead on an NVIDIA A100 GPU using images with a resolution of 448×256.
$[D]$ $[D,M]$ $[D,M]_{e}$
Sec. MB Sec. MB Sec. MB
RIFE [7] 0.03 10.21 0.06 20.46 0.10 30.68
IFRNet [24] 0.14 4.73 0.17 14.91 0.21 25.13
AMT-S [18] 0.20 2.86 0.22 15.01 0.26 25.23
EMA-VFI [17] 2.35 62.62 2.38 74.84 2.42 85.04
VI-I Computational costs of the proposed framework
Distance indexing
Transitioning from time indexing ($[T]$) to distance indexing ($[D]$) does not introduce extra computational costs during the inference phase, yet significantly enhancing image quality. In the training phase, the primary requirement is a one-time offline computation of distance maps for image triplets.
Iterative reference-based estimation
Given that the computational overhead of merely expanding the input channel, while keeping the rest of the structure unchanged, is negligible, the computational burden during the training phase remains equivalent to that of the $[D]$ model. Regarding inference, the total consumption is equal to the number of iterations $\times$ the consumption of the $[D]$ model. We would like to highlight that this iterative strategy is optional: Users can adopt this strategy at will when optimal interpolation results are demanded and the computational budget allows.
Continuous indexing map estimation
The runtime and the number of parameters are reported in Table IX. The estimation process introduces an additional latency of approximately 0.04 seconds per inference, which is relatively low. The number of parameters for the map estimator is 10.2M. Overall, the computational overheads introduced by indexing map estimation are acceptable in practical scenarios.
Multi-frame refiner
We also evaluate the computational overhead introduced by the multi-frame refiner as reported in Table IX. The refiner adds approximately 0.03 seconds per inference and requires an additional 10.3 million parameters, which is practical for deployment.
VII Conclusion and Future Work
@@
We challenge the traditional time indexing paradigm and address its inherent uncertainties related to velocity distribution. Through the introduction of distance indexing and iterative reference-based estimation strategies, we offer a transformative paradigm to VFI. Our innovative plug-and-play strategies not only improves the performance in video interpolation but also empowers users with granular control over interpolation patterns across varied objects. We also propose a continuous distance map estimator to accurately predict distance maps when using multi-frame inputs. Additionally, a multi-frame refiner is integrated into the interpolation pipeline for further enhancement. While the proposed framework significantly improves both pixelwise and perceptual metrics, it still faces challenges in learning and representing diverse interpolation trajectories. Extreme large motions may require generative priors from models like diffusion models. The insights gleaned from our strategies have potential applications across a range of tasks that employ time indexing, such as space-time super-resolution, future predictions, blur interpolation and more.
Acknowledgments
@@
This work was partially supported by the National Key R&D Program of China (No. 2022ZD0160104). We thank Dorian Chan, Zhirong Wu, and Stephen Lin for their insightful feedback and advice. Our thanks also go to Vu An Tran for developing the web application.
Appendix A Proof of Eq. 4
Eq. 4 in the paper can be rigorously proven if an L2 loss is used,
$\min_{\hat{I}_{t}}L=\mathbb{E}_{I_{t}\sim\mathcal{F}(I_{0},I_{1},t)}[(\hat{I}_{t}-I_{t})^{2}].$ (19)
By setting the gradient to zero (this assumes during training, the neural network can reach the exact solution at this point),
$\frac{\partial L}{\partial\hat{I}_{t}}$ $=0$ (20)
$\mathbb{E}_{I_{t}\sim\mathcal{F}(I_{0},I_{1},t)}[\frac{\partial}{\partial\hat{I}_{t}}(\hat{I}_{t}-I_{t})^{2}]$ $=0$
$\mathbb{E}_{I_{t}\sim\mathcal{F}(I_{0},I_{1},t)}[2(\hat{I}_{t}-I_{t})]$ $=0$
$\mathbb{E}_{I_{t}\sim\mathcal{F}(I_{0},I_{1},t)}[\hat{I}_{t}]-\mathbb{E}_{I_{t}\sim\mathcal{F}(I_{0},I_{1},t)}[I_{t}]$ $=0$
$\hat{I}_{t}$ $=\mathbb{E}_{I_{t}\sim\mathcal{F}(I_{0},I_{1},t)}[I_{t}]$
Even when a different loss is used, we still observe a similar phenomenon in practice.
Appendix B The Rationale for Solving Ambiguity
First and foremost, it is essential to clarify that velocity ambiguity can solely exist and be resolved in the training phase, not in the inference phase. The key idea behind our approach can be summarized as follows: While conventional VFI methods with time indexing rely on a one-to-many mapping, our distance indexing learns an approximate one-to-one mapping, which resolves the ambiguity during training. When the input-output relationship is one-to-many during training, the training process fluctuates among conflicting objectives, ultimately preventing convergence towards any specific optimization goal. In VFI, the evidence is the generation of blurry images in the inference phase. Once the ambiguity has been resolved using the new indexing method in the training phase, the model can produce significantly clearer results regardless of the inference strategy used.
Indeed, this one-to-many ambiguity in training is not unique to VFI, but for a wide range of machine learning problems. It is sometimes referred to as “mode averaging” in the community. [^1] https://www.cs.toronto.edu/~hinton/coursera/lecture13/lec13.pdf In some areas, researchers have come up with similar methods [52, 60].
A specific instantiation of this problem
Let us look at an example in text-to-speech (TTS). The same text can be paired with a variety of speeches, and direct training without addressing ambiguities can result in a “blurred” voice (a statistical average voice). To mitigate this, a common approach is to incorporate a speaker embedding vector or a style embedding vector (representing different gender, accents, speaking styles, etc.) during training, which helps reduce ambiguity. During the inference phase, utilizing an average user embedding vector can yield high-quality speech output. Furthermore, by manipulating the speaker embedding vector, effects such as altering the accent and pitch can also be achieved.
Here is a snippet from a high-impact paper Wang et al. [52] which came up with the style embedding in TTS:
Many TTS models, including recent end-to-end systems, only learn an averaged prosodic distribution over their input data, generating less expressive speech – especially for long-form phrases. Furthermore, they often lack the ability to control the expression with which speech is synthesized.
Understanding this example can significantly help understand our paper, as there are many similarities between the two, e.g., motivation, solution, and manipulation.
A minimal symbolic example to help understand better
Assuming we want to train a mapping function $\mathcal{F}$ from numbers to characters.
Training input-output pairs with ambiguity ($\mathcal{F}$ is optimized):
$1\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}a,1\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}b,2\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}a,2\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}b$
$\mathcal{F}$ is optimized with some losses involving the input-output pairs above:
$\underset{\mathcal{F}}{\min}\quad L(\mathcal{F}(1),a)+L(\mathcal{F}(1),b)+L(\mathcal{F}(2),a)+L(\mathcal{F}(2),b),$
where $L$ can be L1, L2 or any other kind of losses. Because the same input is paired with multiple different outputs, the model $\mathcal{F}$ is optimized to learn an average (or, generally, a mixture) of the conflicting outputs, which results in blur at inference.
Inference phase ($\mathcal{F}$ is fixed):
$1\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}\{a,b\}?,2\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}\{a,b\}?$
Training without ambiguity ($\mathcal{F}$ is optimized):
$1\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}a,1\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}a,2\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}b,2\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}b$
In the input-output pairs above, each input value is paired with exactly one output value. Therefore, $\mathcal{F}$ is trained to learn a unique and deterministic mapping.
Inference phase ($\mathcal{F}$ is fixed):
$1\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}a,2\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}b$
Coming back to VFI
When time indexing is used, the same $t$ value is paired to images where the objects are located at various locations due to the speed and directional ambiguities. When distance mapping is used, a single $d$ value is paired to images where the objects are always at the same distance ratio, which allows the model to learn a more deterministic mapping for resolving the speed ambiguity.
It is important to note that fixing the ambiguity does not solve all the problems: At inference time, the “correct” (close to ground-truth) distance map is not available. In this work, we show that it is possible to provide uniform distance maps as inputs to generate a clear output video, which is not perfectly pixel-wise aligned with the ground truth. This is the reason why the proposed method does not achieve state-of-the-art in terms of PSNR and SSIM in Tab. 1 of the paper. However, it achieves sharper frames with higher perceptual quality, which is shown by the better LPIPS and NIQE.
We claim the “correct” distance map is hard to estimate accurately from merely two frames since there are a wide range of possible velocities. If considering more neighboring frames like Xu et al. [8] (more observation information), it is possible to estimate an accurate distance map for pixel-wise aligned interpolation, which we leave for future work.
Furthermore, manipulating distance maps corresponds to sampling other possible unseen velocities, i.e., 2D manipulation of frame interpolation, similar to that mentioned TTS paper Wang et al. [52].
Figure 11: Additional comparison of qualitative results. [T] RIFE (Tri) denotes RIFE [7] trained in a fixed time indexing paradigm (Vimeo90K triplet dataset [54]). [D] RIFE (Tri) denotes the model trained using distance indexing. All models use uniform maps.
TABLE X: Comparison on Vimeo90K triplet dataset. $[T]$ denotes the method trained with traditional fixed time indexing paradigm. $[D]$ denotes the distance indexing paradigm. $[\cdot]_{u}$ denotes inference with uniform map as time indexes.
RIFE [7] IFRNet [24] AMT-S [18] EMA-VFI [17]
$[T]$ $[D]$ $[D]_{u}$ $[T]$ $[D]$ $[D]_{u}$ $[T]$ $[D]$ $[D]_{u}$ $[T]$ $[D]$ $[D]_{u}$
[…]
36.04
35.18 35.80
36.26
35.14 35.97
36.56
35.21 36.50
37.13
36.21
[…]
0.979
0.976 0.979
0.981
0.977 0.980
0.982
0.977 0.982
0.983
0.981
[…]
0.022
0.023 0.020
0.019
0.021 0.021
0.020
0.023 0.020
0.019
0.020
[…]
5.224
5.256 5.245
5.225
5.308 5.293
5.288
5.372 5.343
5.335
Appendix C Additional Experiments and Analysis
Comparison of the fixed-time setting
While the benefits of our proposed disambiguation strategies are best demonstrated on arbitrary-time VFI models, they actually improve the performance of fixed-time models as well. Using RIFE [7] as a representative example, we extend our comparison to the fixed-time training paradigm, depicted in Fig. 11. The label $[T]$ RIFE (Tri) refers to the model trained on the triplet dataset from Vimeo90K [54] employing time indexing. Conversely, $[D]$ RIFE (Tri) indicates training on the same triplet dataset but utilizing our distance indexing approach. Both $[D]$ RIFE and $[D,R]$ RIFE models are trained on the septuplet dataset, consistent with our earlier comparison. Despite being trained on varied datasets, it is evident that the arbitrary time model outperforms the fixed time model. However, the efficacy of distance indexing appears restrained within the fixed-time training paradigm. This limitation stems from the fact that deriving distance representation solely from the middle frame yields a sparse distribution, making it challenging for the network to grasp the nuances of distance. We delve deeper into the quantitative analysis of these findings in Table X. Compared to training arbitrary time models on the septuplet dataset, the advantages of distance indexing become notably decreased when training fixed time models on the triplet dataset.
TABLE XI: Comparison on the septuplet of Vimeo90K [54] using LPIPS loss [56]. We use RIFE [7] as a representative example.
$[T]$ $[D]_{u}$ $[D,R]_{u}$
PSNR $\uparrow$
27.19
26.71 26.72
SSIM $\uparrow$
0.898
0.889 0.890
LPIPS $\downarrow$
0.061
0.065 0.064
NIQE $\downarrow$ 6.307 5.901
5.837
Comparison of using perceptual loss
In addition to training with traditional pixel losses based on L1 and L2 losses, we present the results of employing the more recent LPIPS loss [56] with a VGG backbone [61], as shown in Table XI. The non-reference perceptual quality metric, NIQE, shows notable improvement across all variants. The results also consistently demonstrate the effectiveness of our strategies in resolving velocity ambiguity. Besides, due to the direct optimization of LPIPS loss, the assumption of constant speed in uniform maps affects the performance for this metric. This is why in the test results, $[T]$ has a lower LPIPS, while $[D]_{u}$ and $[D,R]_{u}$ are slightly higher.
TABLE XII: Comparison with $[D_{x,y}]$ scheme on Vimeo90K septuplet dataset. $[D_{x,y}]$ denotes training with two channel distance map with $\bm{x}$ and $\bm{y}$ directions. $[\cdot]_{u}$ denotes inference with uniform maps.
$[T]$ $[D]_{u}$ $[D,R]_{u}$ $[D_{x,y}]_{u}$ $[D_{x,y},R]_{u}$
[…]
0.086
0.091 0.087
[…]
6.220
6.296
6.220
Figure 12: User interface of user study.
Two channel scheme
Comparison results with the two channel scheme, i.e., $D_{t}$ with $\bm{x}$ and $\bm{y}$ directions, are shown in Table XII. The observations are as follows: (1) $[D_{x,y}]_{u}$ outperforms $[D]_{u}$, which makes sense since it accounts for both speed and direction. (2) $[D,R]_{u}$ performs better than $[D_{x,y}]_{u}$. Our speculation is that $[R]$ benefits not only from addressing the directional ambiguity but also from reducing the prediction difficulty via divide-and-conquer. (3) $[D_{x,y},R]_{u}$ does not exceed $[D,R]_{u}$, showing that the iterative formulation is sufficient to resolve the directional ambiguity. $[D_{x,y}]_{u}$ has the potential to learn different trajectories but cannot realize that potential since the trajectory distribution within a short time is not sufficiently diverse. However, $[D_{x,y}]_{u}$ only involves increasing the input channels and does not need to run iteratively. Thus, $[D_{x,y}]_{u}$ is a better choice for fast/lightweight frame interpolation.
Why did $[R]$ fail on AMT-S?
As compared to the speed ambiguity, the directional ambiguity only has a minor impact to the interpolation quality due to the short time span between frames. Merely employing iterative reference-based estimation without tackling speed ambiguity $[T,R]$ can result in the accumulation of inaccuracies. This phenomenon is especially evident in AMT-S, which we attribute to its scaled lookup operation of bidirectional 4D correlation volumes. Cumulative errors are exacerbated in inaccurate iterative lookup operations.
Why NIQE performs better with uniform maps?
Uniform maps tend to yield better results due to smoothing. Nonuniform maps consist of unavoidable inaccuracies from flow estimation, which may introduce unnatural details, resulting in worse NIQE scores but are not noticeable to human perception.
Appendix D User Study UI
As shown in Fig. 12, we initially presented users with the input starting and ending frames. Subsequently, the results from each model’s four distinct variants were displayed anonymously in a sequence, with the order shuffled for each presentation. Users were tasked with reordering the images by dragging them, placing them from left to right based on their perceived quality, i.e., the best image on the extreme left and the least preferred on the far right.
Appendix E Demo
We have included a video demo (available in supplementary materials named “supp.mp4”) to intuitively showcase the enhanced quality achieved through our strategies. The video further illustrates the idea of manipulating object interpolations and provides a guide on using the related web application.